OpenStudy (anonymous):

If the APR of a savings account is 5.3% and interest is compounded quarterly, what is the approximate APY of the account?

OpenStudy (wolf1728):

Go to this page http://www.1728.org/rate.htm and use the first formula: Compounded Rate = (1 + (r/n)^n) -1 Compounded Rate = (1 + (.053/4)^4) -1 Compounded Rate = 1.0540627106 -1 Compounded Rate = 5.40627106%